OpenStudy (anonymous):

Surds help? Solve for a and b. The square root of (a plus the (square root of 3 times b)) is equal to 13 all over 4 plus the square root of 3. Thanks in advance! :)

OpenStudy (anonymous):

\[\sqrt{a+b \sqrt{3}} = \frac{13}{4+\sqrt{3}}\]

OpenStudy (anonymous):

Already figured out that the left hand side is 4 - sqrt of 3.

OpenStudy (anonymous):

When you get finished in here could you please have a look at mine

OpenStudy (anonymous):

square both sides first and then rationalise the denominator in the RHS.And then equate the like terms.

OpenStudy (anonymous):

Alright, Yahad. I just seem to be getting different answers every time is all. :) Thank you.